OpenStudy (anonymous):

Okay so this portfolio tells me to use matrices to find the combined amount of elements found in soil for each sample. How do I do that? is it just normal matrix multiplication or do I have to do something new?

OpenStudy (anonymous):

This is in regards to activity 2.

OpenStudy (e.mccormick):

\( \begin{bmatrix} a&b&c \end{bmatrix} + \begin{bmatrix} d&e&f \end{bmatrix} = \begin{bmatrix} a+d&b+e&c+f \end{bmatrix} \)

OpenStudy (e.mccormick):

So you add up all 5 matrices.

OpenStudy (e.mccormick):

Or all 4 matrices, I mean, to get all 5 answers.

OpenStudy (anonymous):

im stuck on this too ! :'(

OpenStudy (anonymous):

How do i get 5 answers? @e.mccormick

OpenStudy (anonymous):

Oh never mind! Sorry finals have me losing my mind. The answer would be [19.4 14.23 54.6 21.5 4.5] correct?

OpenStudy (e.mccormick):

I don't see hot the last one would be that low with the 26 xylene.